OpenStudy (anonymous):

I am reading that Euclid's algorithm can give you two integers, S and T such that the greatest common factor for two integers N and M gcf(M,N) = SM + TN ... but this seems mathematically impossible. Is this a typo in the book? How does this work? Example, gcf(12,16) is 4. Simple. But there are no positive integers S and T such that 12S + 16T = 4. What am I missing?

OpenStudy (loser66):

You can apply matrix method to find it out |1 0 16| | 1 -1 4| | 1 -1 4| | 0 1 12| -R2 +R1 --> | 0 1 12| and -3*R1 +R2 | -3 4 0| which gives us 4 = 1(12) + (-1) 16 therefore, S =1 and T =-1 We can't have both them are positive.

OpenStudy (loser66):

Other way is algebraic way gcd( 12, 16) 16 = 1(12) +4 12 = 3(4) +0 so gcd (12,16) =4